OpenStudy (anonymous):

Five times the sum of the digits of a two-digit number is 13 less than the original number. If you reverse the digits in the two-digit number, four times the sum of its two digits is 21 less than the reversed two-digit number. The difference of the original two-digit number and the number with reversed digits is

OpenStudy (shinalcantara):

Let 'x' be the number in the tens place 'y' be the number in the units place Then the original number would be x(10) + y 10x + y ----original two-digit number 10y + x ----reversed two-digit number --------------------------------- Five times the sum of the digits of a two-digit number is 13 less than the original number 5(x+y) = (10x + y) - 13 --------------------------------- Distribute 5 to the quantity (x+y) and remove parenthesis 5x + 5y = 10x + y - 13 --------------------------------- Transpose 10x and y to the left side of the equation 5x - 10x + 5y - y = -13 -------------------------------- Combine like terms -5x + 4y = -13 ------equation 1 -------------------------------- If you reverse the digits in the two-digit number, four times the sum of its two digits is 21 less than the reversed two-digit number 4(x+y) = 10y + x -21 -------------------------------- Distribute 4 to the quantity (x+y) and remove parenthesis 4x + 4y = 10y + x - 21 ------------------------------- Transpose 10y and x to the left side of the equation 4x - x + 4y - 10y = -21 ------------------------------- Combine like terms 3x - 6y = -21 ------equation 2 ------------------------------- Multiply equation 1 with 3 and equation 2 with 5 (-5x + 4y = -13) x3 -15x + 12y = -39 ------equation 1' ---------------- (3x - 6y = -21) x5 15x - 30y = -105 -----equation 2' ---------------- Add equation 1' and equation 2' -15x + 12y = -39 + 15x - 30y = -105 -18y = -144 y = 8 ------------------------------ Substitute y=8 to equation 1 -5x + 4y = -13 -5x + 4(8) = -13 -5x + 32 = -13 Transpose 32 to the right side -5x = -13 - 32 -5x = -45 x = 9 ----------------------------- 10x + y -----original number 10y + x -----reversed number ----------------------------- The difference would be 10x + y - (10y + x) = [10(9) + 8] -[10(8)+ 9] = 98 - 89 = 9 ----------------------------- Therefore the difference between the original and the reversed is 9.